Using the marginal income tax rates shown below, what is the average tax rate on incomes of $100,000 and $200,000? Give all answers to two decimals.



a. Average tax rate on $100,000 of income:    %

b. Average tax rate on $200,000 of income:    %
Answer #1

Here, 

Taxable incomeTax Rate
0 - 907510%
9076 - 3690015%
36901 - 8935025%
89351 - 18635028%
186351 - 40510033%
405101 - 40675035%
Over 40675139.6%

Now for $100000, tax will be

= 10%* 9075 + 15%* (36900-9075)+25%*(89350-36900) + 28%*(100000 - 89350)

= 907.5 + 4173.75 + 13112.5+2981.72 = 21175.47

So Average Tax rate will be = 21175.47/ 100000 = 21.18%

Now for $200000, tax will be

= 10%* 9075 + 15%* (36900-9075)+25%*(89350-36900) + 28%*(186350 - 89350)+33%*(200000 - 186350)

= 907.5 + 4173.75 + 13112.5+27160+4504.5 = 49858.25

So Average Tax rate will be = 49858.25/ 200000 = 24.93%


answered by: sunraise